OpenStudy (anonymous):

Using trigonometric identities, transform the left side of the equation into the right side: (1+cos(x))(1-cos(x)) = sin^2(x)

OpenStudy (anonymous):

Expand the left hand side.

OpenStudy (anonymous):

Multiply it out.

OpenStudy (anonymous):

Do you know how to multiply them?

OpenStudy (anonymous):

I used the FOIL method and got cos^4(x) = sin^2(x).

OpenStudy (anonymous):

You did not foil responsibly.

OpenStudy (anonymous):

\[\begin{split} (1+\cos(x))(1-\cos(x)) &=(1-\cos(x))+\cos(x)(1-\cos(x)) \\ &=1-\cos(x)+\cos(x)-\cos(x)\cos(x) \\ &=1-\cos^2(x) \end{split} \]